Can somebody help me solve this problem?
Solve: x-2/5 x y=-2
3x+2y = 0
solve for x and y.

Answer by Alan3354(69443)

Can somebody help me solve this problem?
Solve: x-2/5 x y=-2
3x+2y = 0
solve for x and y.
-------------------
Is there an xy term?
x - (2/5xy) = -2 or
(x-2)/5xy = -2 ?